# MAD-UV: The 1st INTERSPEECH Mice Autism Detection via Ultrasound Vocalization Challenge

### v1

@@ Abstract
The Mice Autism Detection via Ultrasound Vocalization (MAD-UV) Challenge introduces the first INTERSPEECH challenge focused on detecting autism spectrum disorder (ASD) in mice through their vocalizations. Participants are tasked with developing models to automatically classify mice as either wild-type or ASD models based on recordings with a high sampling rate. Our baseline system employs a simple CNN-based classification using three different spectrogram features. Results demonstrate the feasibility of automated ASD detection, with the considered audible-range features achieving the best performance (UAR of $0.600$ for segment-level and $0.625$ for subject-level classification). This challenge bridges speech technology and biomedical research, offering opportunities to advance our understanding of ASD models through machine learning approaches. The findings suggest promising directions for vocalization analysis and highlight the potential value of audible and ultrasound vocalizations in ASD detection.
@@ 1 Introduction
One area that has garnered interest in this context is the analysis of Ultrasound Vocalizations (USVs) produced by mice [9]. These vocalizations, inaudible to humans, have been shown to vary significantly between wild-type (healthy) mice and those genetically engineered to model ASD [4]. In this work, we introduce the Mice Autism Detection via Ultrasound Vocalization Challenge (MAD-UV), in which participants are tasked with building models to automatically classify mice as either ASD or wild-type based on their USVs. The challenge makes use of data collected in the study [4], comprising around $7$ hours of USVs by $84$ different subjects. While numerous previous studies have analyzed human vocalizations for ASD detection [10, 11, 12], MAD-UV extends the scope to non-human subjects, making it a unique problem for computational models to solve. Given the plethora of works addressing different aspects of human speech, participants are particularly encouraged to explore the potential of transferring models and techniques on human data to ultrasonic vocalizations from mice. The baseline experiments for MAD-UV as presented in this paper demonstrate the feasibility of the general approaches. The potential of leveraging methods that were proven in human speech for detecting autism in mice is indicated by the baseline results presented in this paper, highlighting the versatility of machine learning and its ability to bridge gaps between species in behavioral and neurological research. Thus, as the first animal health challenge in INTERSPEECH, MAD-UV provides an opportunity to contribute to both speech technology and biomedical research by advancing our understanding of ASD in different species through machine learning for speech data.
@@ 2 Related Work
Though mice also produce sounds in a range audible to humans, their ultrasound vocalizations are more frequent and have been proven informative with regard to a range of characteristics, including sex, age, and health conditions [22]. Ivanenko et al. [23] successfully utilized deep neural networks to determine the sex of mice based on USVs. Vogel et al. [24] used supervised learning with random forests to predict $9$ expert-defined mice vocalization types, achieving $85$ classification accuracy. An alternative to expert-defined categories, Wang et al. [25] propose an unsupervised spectral clustering approach to identify mice USV types. In the most similar work to MAD-UV, Qian et al. [26] conducted a pilot study using a large-scale pre-trained audio neural network to extract high-level features from mice USVs for identifying ASD model mice. Their approach achieved a UAR of $66.6$ in this binary classification task, marking the first use of inaudible USVs for detecting ASD type in mice.
@@ 3 Data
For the training and validation sets, participants will have access to the complete $5$-minute audio recordings, allowing them free implementation of customized segmentation approaches and training-validation partition strategies. As mentioned above, the test set recordings underwent systematic segmentation into $160$ $15$-second clips. To maintain assessment integrity, all labels of the test set were removed, and the clips were systematically shuffled according to a randomly generated rule that will be used in the evaluation. Comprehensive partition statistics are detailed in Table 1.
@@ 4 Baseline Experiments
MAD-UV introduces two evaluation criteria: segment-level and subject-level classification. During both the training and testing phases, the trained model predicts the class of all segments in the validation and test sets, respectively. To determine the subject-level classification, majority voting is applied to the segment predictions belonging to the same subject. For instance, if a greater number of segments for a subject are classified as wild-type rather than ASD, the subject is categorized as wild-type.
@@ 5 Results
6 Challenge Organization and Outcome
In order to participate in MAD-UV, teams must register under the lead of a professor in academia, or a research team leader in industry. Upon signing the EULA, they obtain access to the dataset, the baseline code as well as the evaluation system. Teams are allowed to comprise at most $5$ members excluding the PI. Moreover, we do not allow one and the same person to be part of several teams. None of the organizers will be members of any participating team.
The evaluation system will be hosted on eval.ai [^3] https://eval.ai/web/challenges/challenge-page/2428/, an established platform for shared task evaluation. This way, there will be a transparent leaderboard and automatic enforcement of the constraints on the number of submissions. Participants submit their segment-level predictions for the test set and immediately receive their results. Note that the test subject IDs are not disclosed to the participants. The subject-level results are calculated automatically using the same majority voting approach as described above.
Analogously to the baseline evaluation (cf. Table 2), we evaluate participants’ submissions on both the segment and the subject level via the respective UARs. This leads to two rankings, which are averaged to obtain an overall rank. In case the two teams’ average ranks are equal, priority is given to the team with a lower rank on the segment level. Officially winning MAD-UV requires to submit a system paper describing the technical approach, which must also be accepted into the conference.
In order to facilitate the reproducibility of the results reported above and enable participants to rapidly get started with developing their approaches, we make the code for our baseline systems publicly available [^4] https://github.com/KamijouMikoto/MAD-UV_2025/. Furthermore, participants can download all $5$ pre-extracted feature sets and the relevant checkpoints of our baseline models. All up-to-date information on the challenge logistics can be found at the MAD-UV website [^5] https://www.mad-uv.org/.
@@ 7 Conclusion
In this baseline paper, we introduced MAD-UV, the 1st INTERSPEECH Mice Autism Detection via Ultrasound Vocalization Challenge. We described the extensive challenge dataset and reported a set of competitive baseline results that serve as benchmarks for participants’ approaches. While all feature sets considered lead to above-chance performance, the results obtained with the audible spectrogram prove to be particularly encouraging. These insights not only enhance our understanding of mice vocalizations but also pave the way for leveraging human speech technologies to better analyze animal communication and behaviors. MAD-UV is designed to enable participants to contribute to this intriguing field within a standardized shared task setup. Besides those aiming at the merely quantitative task of outperforming the baselines, we also look forward to contributions that deal with the dataset in a more qualitative manner and thus increase our understanding of the data.
@@ 8 Acknowledgements
The challenge is supported by the Japan Science and Technology Agency MOONSHOT Program (Grant JPMJMS229B) and the DFG project No. 512414116 (HearTheSpecies). Furthermore, this work was supported by MDSI – the Munich Data Science Institute as well as MCML – the Munich Center of Machine Learning. Björn W. Schuller is also with the Konrad Zuse School of Excellence in Reliable AI (relAI), Munich, Germany.

### v3

# MADUV: The 1st INTERSPEECH Mice Autism Detection via Ultrasound Vocalization Challenge

@@ Abstract
The Mice Autism Detection via Ultrasound Vocalization (MADUV) Challenge introduces the first INTERSPEECH challenge focused on detecting Autism Spectrum Disorder (ASD) in mice through their vocalizations. Participants are tasked with developing models to automatically classify mice as either wild-type or ASD models based on recordings with a high sampling rate. Our baseline system employs a simple CNN-based model using three different features. Results demonstrate the feasibility of automated ASD detection, with the considered audible-range features achieving the best performance (UAR of $0.600$ for segment-level and $0.625$ for subject-level classification). This challenge bridges speech technology and biomedical research, offering opportunities to advance our understanding of ASD models through machine learning approaches. The findings suggest promising directions for vocalization analysis and highlight the potential value of audible and ultrasound vocalizations in ASD detection.
@@ 1 Introduction
One area that has garnered interest in this context is the analysis of Ultrasound Vocalizations (USVs) produced by mice [9]. These vocalizations, inaudible to humans, have been shown to vary significantly between wild-type (healthy) mice and those genetically engineered to model ASD [4]. In this work, we introduce the Mice Autism Detection via Ultrasound Vocalization Challenge (MADUV), in which participants are tasked with building models to automatically classify mice as either ASD or wild-type based on their USVs. The challenge makes use of data collected in the study [4], comprising around $7$ hours of USVs by $84$ different subjects. While numerous previous studies have analyzed human vocalizations for ASD detection [10, 11, 12], MADUV extends the scope to non-human subjects, making it a unique problem for computational models to solve. Given the plethora of works addressing different aspects of human speech, participants are particularly encouraged to explore the potential of transferring models and techniques on human data to ultrasonic vocalizations from mice. The baseline experiments for MADUV as presented in this paper demonstrate the feasibility of the general approaches. The potential of leveraging methods that were proven in human speech for detecting autism in mice is indicated by the baseline results presented in this paper, highlighting the versatility of machine learning and its ability to bridge gaps between species in behavioral and neurological research. Thus, as the first animal health challenge in INTERSPEECH, MADUV provides an opportunity to contribute to both speech technology and biomedical research by advancing our understanding of ASD in different species through machine learning for speech data.
@@ 2 Related Work
Though mice also produce sounds in a range audible to humans, their ultrasound vocalizations are more frequent and have been proven informative with regard to a range of characteristics, including sex, age, and health conditions [22]. Ivanenko et al. [23] successfully utilized deep neural networks to determine the sex of mice based on USVs. Vogel et al. [24] used supervised learning with random forests to predict $9$ expert-defined mice vocalization types, achieving $85$ classification accuracy. An alternative to expert-defined categories, Wang et al. [25] propose an unsupervised spectral clustering approach to identify mice USV types. In the most similar work to MADUV, Qian et al. [26] conducted a pilot study using a large-scale pre-trained audio neural network to extract high-level features from mice USVs for identifying ASD model mice. Their approach achieved a UAR of $66.6$ in this binary classification task, marking the first use of inaudible USVs for detecting ASD type in mice.
@@ 3 Data
For the training and validation sets, participants will have access to the complete $5$-minute audio recordings, allowing them free implementation of customized segmentation approaches and training-validation partition strategies. As mentioned above, the test set recordings underwent systematic segmentation into $160$ $30$-second clips. To maintain assessment integrity, all labels of the test set were removed, and the clips were systematically shuffled according to a randomly generated rule that will be used in the evaluation. The partition statistics are detailed in Table 1.
@@
It should be clarified that this challenge uses mice models to study basic neural development mechanisms, adhering to ethical standards for animal research. The research focuses solely on analysing mice vocalisations and does not aim for human diagnosis or intervention. The challenge contributes to neuroscience by advancing understanding of neural development, while recognising the limitations of animal models in representing human conditions like autism.
@@ 4 Baseline Experiments
MADUV introduces two evaluation criteria: segment-level and subject-level classification. During both the training and testing phases, the trained model predicts the class of all segments in the validation and test sets, respectively. To determine the subject-level classification, majority voting is applied to the segment predictions belonging to the same subject. For instance, if a greater number of segments for a subject are classified as wild-type rather than ASD, the subject is categorized as wild-type.
@@ 5 Results
6 Challenge Organization
In order to participate in MADUV, teams must register under the lead of a professor in academia, or a research team leader in industry. Upon signing the EULA, they obtain access to the dataset, the baseline code as well as the evaluation system. Teams are allowed to comprise at most $5$ members excluding the PI. Moreover, we do not allow one and the same person to be part of several teams. None of the organizers will be members of any participating team.
The evaluation system will be hosted on EvalAI [^3] https://eval.ai/web/challenges/challenge-page/2428/, an established platform for shared task evaluation. This way, there will be a transparent leaderboard and automatic enforcement of the constraints on the number of submissions. Participants submit their segment-level predictions for the test set and immediately receive their results. Note that the test subject IDs are not disclosed to the participants. The subject-level results are calculated using the same majority voting approach above.
Analogously to the baseline evaluation (cf. Table 2), we evaluate participants’ submissions on both the segment and the subject level via the respective UARs. This leads to two rankings, which are averaged to obtain an overall rank. In case the two teams’ average ranks are equal, priority is given to the team with a lower rank on the segment level. Officially winning MADUV requires to submit a system paper describing the technical approach, which must also be accepted into the conference.
In order to facilitate the reproducibility of the results reported above and enable participants to rapidly get started with developing their approaches, we make the code for our baseline systems publicly available [^4] https://github.com/KamijouMikoto/maduv_2025/. Furthermore, participants can download all $5$ pre-extracted feature sets and the relevant checkpoints of our baseline models. All up-to-date information on the challenge logistics can be found at the MADUV website [^5] https://www.maduv.org/.
@@ 7 Conclusion
In this baseline paper, we introduced MADUV, the 1st INTERSPEECH Mice Autism Detection via Ultrasound Vocalization Challenge. We described the extensive challenge dataset and reported a set of competitive baseline results that serve as benchmarks for participants’ approaches. While all feature sets considered lead to above-chance performance, the results obtained with the audible spectrogram prove to be particularly encouraging. These insights not only enhance our understanding of mice vocalizations but also pave the way for leveraging human speech technologies to better analyze animal communication and behaviors. Besides those aiming at the merely quantitative task of outperforming the baselines, we also look forward to contributions that deal with the dataset in a more qualitative manner and thus increase our understanding of the data.
@@ 8 Acknowledgements
The paper is supported by the Japan Science and Technology (JST) Agency MOONSHOT R&D (Grant JPMJMS229B and JPMJMS2021). Furthermore, this paper is supported by MDSI – the Munich Data Science Institute as well as MCML – the Munich Center of Machine Learning. Björn W. Schuller is also with the Konrad Zuse School of Excellence in Reliable AI (relAI), Munich, Germany.